OpenStudy (anonymous):

Which equation will result if the substitution method is used to solve the following system of linear equations? x + y = 6 x = y - 2 (y - 2) + y = 6 x + y = y - 2 x + (y - 2) = 6 x = (x - 6) - 2

OpenStudy (anonymous):

If x is substituted in, you would get (y-2) + y = 6.

OpenStudy (anonymous):

x + y = 6 --- (1) x = y - 2 ----( 2) Substituting (2) in (1): (y - 2) + y = 6

OpenStudy (anonymous):

x+y=6 x-y=-2 2x =4 x =2 y =4 (2,4)